# centertableaux

### v1

@@ 2.4. Counting formula
For $\widetilde{A}$, the above theorem yields the following formula immediately.
@@ 3.1. Standard representations of classical groups
The stabilizer of the flag in $\mathfrak{g}$ is a Borel subalgebra of $\mathfrak{g}$. Using this Borel subalgebra, we get canonical identifications
@@
This identification is independent of the choice of flag (3.1). As in the Section 2.1, we have a canonical identification:
@@ 3.2.1. Integral case
Also use Proposition 2.10, 2.11, and induction by step, we can see
@@ 3.2.2. General case
$\Gamma_{i}:$ $\Lambda_{i}$$\longrightarrow$ $\quad{\mathrm{Irr}}(H_{i})$
@@
$\varphi:$ ${\mathcal{P}}_{\Lambda_{i}}({\mathrm{G}}_{A^{{\mathbb{R}}}}(V_{1}))\times\cdots\times{\mathcal{P}}_{\Lambda_{r}}({\mathrm{G}}_{A^{{\mathbb{R}}}}(V_{r}))$$\longrightarrow$ $\quad{\mathcal{P}}_{\Lambda}(G)$
@@
The map $\varphi$ defined above is $W(\Lambda)$-equivariant and bijective.
@@
For the bijection, note that for every triple $(H,\xi,\Gamma)\in\mathscr{P}_{\Lambda}(G)$, $H$ can be embedding into the Levi subgroup ${\mathrm{G}}_{A^{{\mathbb{R}}}}(V_{1})\times\cdots\times{\mathrm{G}}_{A^{{\mathbb{R}}}}(V_{r})$ after conjugation. The rest follows from direct verification. ∎
We obtain the following isomorphism using this bijection of parameters and Theorem 2.3.
@@
There is an isomorphism of $W(\Lambda)$-representations
(3.11) $\mathrm{Coh}_{\Lambda}({\mathcal{K}}(G))\cong\mathrm{Coh}_{\Lambda_{1}}({\mathcal{K}}({\mathrm{G}}_{A^{{\mathbb{R}}}}(V_{1})))\boxtimes\cdots\boxtimes\mathrm{Coh}_{\Lambda_{r}}({\mathcal{K}}({\mathrm{G}}_{A^{{\mathbb{R}}}}(V_{r}))).$
@@ 3.3.2. General case
$\varphi:$ ${\mathcal{P}}_{\Lambda_{1}}({\mathrm{G}}_{A^{{\mathbb{H}}}}(V_{i}))\times\cdots\times{\mathcal{P}}_{\Lambda_{r}}({\mathrm{G}}_{A^{{\mathbb{H}}}}(V_{r}))$$\longrightarrow$ $\quad{\mathcal{P}}_{\Lambda}(G)$
@@
The map $\varphi$ defined above is $W(\Lambda)$-equivariant and bijective.
@@
There is an isomorphism of $W(\Lambda)$-representations
(3.16) $\mathrm{Coh}_{\Lambda}({\mathcal{K}}(G))\cong\mathrm{Coh}_{\Lambda_{1}}({\mathcal{K}}({\mathrm{G}}_{A^{{\mathbb{H}}}}(V_{1})))\boxtimes\cdots\boxtimes\mathrm{Coh}_{\Lambda_{r}}({\mathcal{K}}({\mathrm{G}}_{A^{{\mathbb{H}}}}(V_{r}))).$
For a nilpotent orbit ${\mathcal{O}}={\mathcal{O}}_{\iota}$, we can get
@@ 3.4.2. General case
$\varphi:$ ${\mathcal{P}}_{\Lambda_{1}}({\mathrm{G}}_{A^{{\mathbb{C}}}}(V_{1}))\times\cdots\times{\mathcal{P}}_{\Lambda_{r}}({\mathrm{G}}_{A^{{\mathbb{C}}}}(V_{r}))$$\longrightarrow$ $\quad{\mathcal{P}}_{\Lambda}(G)$
@@
The map $\varphi$ defined above is $W(\Lambda)$-equivariant and bijective.
@@
There is an isomorphism of $W(\Lambda)$-representations
(3.29) $\mathrm{Coh}_{\Lambda}({\mathcal{K}}(G))\cong\mathrm{Coh}_{\Lambda_{1}}({\mathcal{K}}({\mathrm{G}}_{A^{{\mathbb{C}}}}(V_{1})))\boxtimes\cdots\boxtimes\mathrm{Coh}_{\Lambda_{r}}({\mathcal{K}}({\mathrm{G}}_{A^{{\mathbb{C}}}}(V_{r}))).$
@@ 3.5.2. General case
$\varphi:$ ${\mathcal{P}}_{\Lambda_{0}}({\mathrm{G}}_{A}(W))\times{\mathcal{P}}_{\Lambda_{0}^{\prime}}({\mathrm{G}}_{A}(W^{\prime}))\times{\mathcal{P}}_{\Lambda_{1}}({\mathrm{GL}}(W_{1}))\times\cdots\times{\mathcal{P}}_{\Lambda_{r}}({\mathrm{GL}}(W_{r}))$$\longrightarrow$ $\quad{\mathcal{P}}_{\Lambda}(G)$
@@
$\varphi:$ ${\mathcal{P}}_{\Lambda_{0}}({\mathrm{G}}_{A}(W))\times{\mathcal{P}}_{\Lambda_{0}^{\prime}}^{\prime}({\mathrm{G}}_{\widetilde{A}}(W^{\prime}))\times{\mathcal{P}}_{\Lambda_{1}}({\mathrm{GL}}(W_{1}))\times\cdots\times{\mathcal{P}}_{\Lambda_{r}}({\mathrm{GL}}(W_{r}))$$\longrightarrow$ $\quad{\mathcal{P}}_{\Lambda}(G)$
@@
Then, by a similar argument as Proposition 3.2 and 3.3, we obtain the following result.
@@
The map $\varphi$ defined above is $W(\Lambda)$-equivariant and bijective.
@@
There is an isomorphism of $W(\Lambda)$-representations
$\mathrm{Coh}_{\Lambda}({\mathcal{K}}(G))\cong$ $\mathrm{Coh}_{\Lambda_{0}}({\mathcal{K}}({\mathrm{G}}_{A}(W)))\boxtimes\mathrm{Coh}_{\Lambda_{0}^{\prime}}({\mathcal{K}}({\mathrm{G}}_{A}(W^{\prime})))$
$\boxtimes\mathrm{Coh}_{\Lambda_{1}}({\mathcal{K}}({\mathrm{GL}}(W_{1})))\boxtimes\cdots\boxtimes\mathrm{Coh}_{\Lambda_{r}}({\mathcal{K}}({\mathrm{GL}}(W_{r}))),$
@@
$\mathrm{Coh}_{\Lambda}({\mathcal{K}}(G))\cong$ $\mathrm{Coh}_{\Lambda_{0}}({\mathcal{K}}({\mathrm{G}}_{A}(W)))\boxtimes\mathrm{Coh}_{\Lambda_{0}^{\prime}}({\mathcal{K}}^{\mathrm{gen}}({\mathrm{G}}_{\widetilde{A}}(W^{\prime})))$
$\boxtimes\mathrm{Coh}_{\Lambda_{1}}({\mathcal{K}}({\mathrm{GL}}(W_{1})))\boxtimes\cdots\boxtimes\mathrm{Coh}_{\Lambda_{r}}({\mathcal{K}}({\mathrm{GL}}(W_{r}))),$

### v2

@@ 2.4. Counting formula
For the case $\widetilde{A}$, the above theorem yields the following formula immediately.
@@ 3.1. Standard representations of classical groups
The stabilizer of the flag in $\mathfrak{g}$ is a Borel subalgebra of $\mathfrak{g}$. Using this Borel subalgebra, we get natural identifications
@@
This identification is independent of the choice of flag (3.1). As in the Section 2.1, we have natural identifications:
@@ 3.2.1. Integral case
At the same time, by using Proposition 2.10, 2.11, and induction by step, we obtain
@@ 3.2.2. General case
(3.11) $\Gamma_{i}:$ $\Lambda_{i}$$\longrightarrow$ $\quad{\mathrm{Irr}}(H_{i})$
@@
(3.12) $\varphi:$ ${\mathcal{P}}_{\Lambda_{i}}({\mathrm{G}}_{A^{{\mathbb{R}}}}(V_{1}))\times\cdots\times{\mathcal{P}}_{\Lambda_{r}}({\mathrm{G}}_{A^{{\mathbb{R}}}}(V_{r}))$$\longrightarrow$ $\quad{\mathcal{P}}_{\Lambda}(G)$
@@
The map $\varphi$ defined above is bijective.
@@
Note that for every triple $(H,\xi,\Gamma)\in\mathscr{P}_{\Lambda}(G)$, $H$ can be embedding into the Levi subgroup ${\mathrm{G}}_{A^{{\mathbb{R}}}}(V_{1})\times\cdots\times{\mathrm{G}}_{A^{{\mathbb{R}}}}(V_{r})$ after conjugation. The rest follows from direct verification. ∎
The bijection of parameters above induces a linear map:
(3.13) $\varphi:$ $\mathrm{Coh}_{\Lambda_{1}}({\mathcal{K}}({\mathrm{G}}_{A^{{\mathbb{R}}}}(V_{1})))\otimes\cdots\otimes\mathrm{Coh}_{\Lambda_{r}}({\mathcal{K}}({\mathrm{G}}_{A^{{\mathbb{R}}}}(V_{r})))$$\longrightarrow$ $\quad\mathrm{Coh}_{\Lambda}({\mathcal{K}}(G))$
[…]
@@
The linear map $\varphi$ defined above is an isomorphism of representations of $W_{\Lambda}$.
Proof.
It is straightforward to verify that the bijection (3.12) is also equivariant with respect to the cross action of $W_{\Lambda}$ and the operation known as the Cayley transformation, as defined in [Vog81, Chapter 8]. The proposition then follows from the explicit description of the coherent continuation representation with respect to the standard basis, as given in [Vog82, Definition 14.4]. ∎
@@ 3.3.2. General case
(3.18) $\varphi:$ ${\mathcal{P}}_{\Lambda_{1}}({\mathrm{G}}_{A^{{\mathbb{H}}}}(V_{i}))\times\cdots\times{\mathcal{P}}_{\Lambda_{r}}({\mathrm{G}}_{A^{{\mathbb{H}}}}(V_{r}))$$\longrightarrow$ $\quad{\mathcal{P}}_{\Lambda}(G)$
@@
The map $\varphi$ defined above bijective.
@@
There is a natural isomorphism of $W(\Lambda)$-representations
(3.19) $\varphi:\quad\mathrm{Coh}_{\Lambda_{1}}({\mathcal{K}}({\mathrm{G}}_{A^{{\mathbb{H}}}}(V_{1})))\boxtimes\cdots\boxtimes\mathrm{Coh}_{\Lambda_{r}}({\mathcal{K}}({\mathrm{G}}_{A^{{\mathbb{H}}}}(V_{r})))\longrightarrow\mathrm{Coh}_{\Lambda}({\mathcal{K}}(G)).$
For a nilpotent orbit ${\mathcal{O}}={\mathcal{O}}_{\iota}$, we obtain
@@ 3.4.2. General case
(3.32) $\varphi:$ ${\mathcal{P}}_{\Lambda_{1}}({\mathrm{G}}_{A^{{\mathbb{C}}}}(V_{1}))\times\cdots\times{\mathcal{P}}_{\Lambda_{r}}({\mathrm{G}}_{A^{{\mathbb{C}}}}(V_{r}))$$\longrightarrow$ $\quad{\mathcal{P}}_{\Lambda}(G)$
@@
The map $\varphi$ defined above is bijective.
@@
There is a natural isomorphism of $W(\Lambda)$-representations
(3.33) $\varphi:\quad\mathrm{Coh}_{\Lambda_{1}}({\mathcal{K}}({\mathrm{G}}_{A^{{\mathbb{C}}}}(V_{1})))\boxtimes\cdots\boxtimes\mathrm{Coh}_{\Lambda_{r}}({\mathcal{K}}({\mathrm{G}}_{A^{{\mathbb{C}}}}(V_{r})))\longrightarrow\mathrm{Coh}_{\Lambda}({\mathcal{K}}(G)).$
@@ 3.5.2. General case
(3.46) $\varphi:$ ${\mathcal{P}}_{\Lambda_{0}}({\mathrm{G}}_{A}(W))\times{\mathcal{P}}_{\Lambda_{0}^{\prime}}({\mathrm{G}}_{A}(W^{\prime}))\times{\mathcal{P}}_{\Lambda_{1}}({\mathrm{GL}}(W_{1}))\times\cdots\times{\mathcal{P}}_{\Lambda_{r}}({\mathrm{GL}}(W_{r}))$$\longrightarrow$ $\quad{\mathcal{P}}_{\Lambda}(G)$
@@
(3.47) $\varphi:$ ${\mathcal{P}}_{\Lambda_{0}}({\mathrm{G}}_{A}(W))\times{\mathcal{P}}_{\Lambda_{0}^{\prime}}^{\prime}({\mathrm{G}}_{\widetilde{A}}(W^{\prime}))\times{\mathcal{P}}_{\Lambda_{1}}({\mathrm{GL}}(W_{1}))\times\cdots\times{\mathcal{P}}_{\Lambda_{r}}({\mathrm{GL}}(W_{r}))$$\longrightarrow$ $\quad{\mathcal{P}}_{\Lambda}(G)$
@@
The map $\varphi$ defined above is bijective.
@@
There is a natural isomorphism of $W(\Lambda)$-representations
$\varphi:\quad$ $\mathrm{Coh}_{\Lambda_{0}}({\mathcal{K}}({\mathrm{G}}_{A}(W)))\boxtimes\mathrm{Coh}_{\Lambda_{0}^{\prime}}({\mathcal{K}}({\mathrm{G}}_{A}(W^{\prime})))\boxtimes\mathrm{Coh}_{\Lambda_{1}}({\mathcal{K}}({\mathrm{GL}}(W_{1})))\boxtimes\cdots$
$\boxtimes\mathrm{Coh}_{\Lambda_{r}}({\mathcal{K}}({\mathrm{GL}}(W_{r})))\longrightarrow\mathrm{Coh}_{\Lambda}({\mathcal{K}}(G)),$
@@
$\varphi:\quad$ $\mathrm{Coh}_{\Lambda_{0}}({\mathcal{K}}({\mathrm{G}}_{A}(W)))\boxtimes\mathrm{Coh}_{\Lambda_{0}^{\prime}}({\mathcal{K}}^{\mathrm{gen}}({\mathrm{G}}_{\widetilde{A}}(W^{\prime})))\boxtimes\mathrm{Coh}_{\Lambda_{1}}({\mathcal{K}}({\mathrm{GL}}(W_{1})))\boxtimes\cdots$
$\boxtimes\mathrm{Coh}_{\Lambda_{r}}({\mathcal{K}}({\mathrm{GL}}(W_{r})))\longrightarrow\mathrm{Coh}_{\Lambda}({\mathcal{K}}(G)),$